OpenStudy (anonymous):

What are the factors of 3x2 + 13x + 4?

OpenStudy (anonymous):

1) 8^1/2 = 2^x 8 can be rewritten as 2^3, so (2^3)^1/2, multiply 3 by 1/2 so the answer is 2 ^ 3/2 2) √3 can be rewritten as 3^1/2 so 3^1/2 = 9^x 9 can be rewritten as 3^2 so 3^1/2 = (3^2)^x x= 1/4 so the answer is 9^1/4 3) 1 / (4√2) can be rewritten as 1 / (2^2 * 2^1/2) anything looking like 1/x can be rewritten as x^-1, or something to the negative power 1 / (2^5/2) add the two powers together to get 5/2 and apply the rule above and the answer is = 2^-5/2

OpenStudy (anonymous):

Mark my answer as correct, if it helped. :)

OpenStudy (anonymous):

Click "Best answer". :)

Parth (parthkohli):

Err. I just noticed that you answered something else

Parth (parthkohli):

You have to factor the polynomial by splitting the middle term and grouping.\[3x^2+ 12x+x+4\]